OpenStudy (anonymous):

Simplify 25 – 10 5 + 3 x 2 9 11 17 29

OpenStudy (eujc21):

\[25 - 10 (?) 5 + (3*2)\]

OpenStudy (eujc21):

\[25 - 10(?)5 + 6 = 19 - 10(?)5 \]

OpenStudy (anonymous):

25 – 10 / 5 + 3 x 2

OpenStudy (anonymous):

division sign was missing. sorry

OpenStudy (eujc21):

ok \[ 25 - (10/5) + (3*2) = 25 -2 +6 = .....\]

OpenStudy (eujc21):

You're fine no problem... but that's how you solve it

OpenStudy (anonymous):

so answer ? cause i keep solving ti wrong

OpenStudy (eujc21):

29

OpenStudy (anonymous):

thnk you sir